OpenStudy (anonymous):

I'm slightly confused on how to do this, could someone explain it for me? Here's the question: Use the substitution method to solve the system of equations. Enter your answer as an ordered pair: x+y=12, -x=-y-10

OpenStudy (anonymous):

I would first multiply the last equation by -1. -1(-x)= -1(-y-10) to get x=y+10

OpenStudy (anonymous):

then substitute that into the other equation. y+10+y=12, 2y+10=12, 2y=2, y=1

OpenStudy (anonymous):

Thank you so much!